9/2/16
Summary:
Light winds out of the Northeast. All stations coming in.
we plan to visit station 15 to work on the rad/soil mote issue. Also visit 2,3,4 to add power/monitoring cable.
Actions past 24 hours:
To dos:
Sensor status:
T/RH: ok
P: ok
2D Gill: ok, some spiking at sites 8,9,10
csat u,v: ok
csat ldiag: spiking up to 1
csat w, tc: ok
EC150: ok
motes: ID2 not giving soil data, ID1 has issues with message sent
Wetness: unknown
radiation: unknown
Tsoil: unknown
Gsoil: unknown
Qsoil: unknown
Cvsoil: unknown
Battery: ok except 2,3,4,14 missing cable